The age of a dad in years is three times the sum of his two children's ages. In four years time the dad is going to be two times older than the sum of his children's ages. How old is the dad?
Answer by ankor@dixie-net.com(22740)

This problem ignores reality, however, lets plow thru it anyway
let d = dad's present age
Let s = sum of the children's present age
:
The age of a dad in years is three times the sum of his two children's ages.
d = 3s
:
In four years time the dad is going to be two times older than the sum of his children's ages.
(d + 4) = 2(s + 4)
d + 4 = 2s + 8
d = 2s + 8 - 4
d = 2s + 4
:
How old is the dad?
:
Substitute 3s for d in the above equation, find s
3s = 2s + 4
3s - 2s = 4
s = 4, sum of the children's ages
:
Then dad is 12!!